# Wave–Particle Complementarity as the Optimal Limit of Unambiguous Quantum-State Discrimination

### v2

@@ III Experimental Protocol
Proposed experimental implementation.— The IDP measurement can be implemented [Barnett, mohseni] directly on the idler modes of a two-crystal ZWM interferometer [Torres2024]. Two coherently pumped SPDC crystals (A,B) generate signal–idler pairs in the low-gain regime. The idler from A traverses an object $S$ with transmissivity $(t,r)$, while the idler from B bypasses it. The signal modes are combined at a balanced beam splitter to measure the single-photon visibility $V$.
@@
To realize the IDP POVM $\{\Pi_{0},\Pi_{1},\Pi_{2}\}$ on the idler, detector $D_{1}$ implements the projection onto $|1,0\rangle$, yielding conclusive outcome $\Pi_{1}$. Detector $D_{2}$ receives the output of a variable beam splitter that interferes the idler modes with a calibrated phase such that the detected mode is $t^{*}|1,0\rangle-r^{*}|0,1\rangle$, orthogonal to $|\phi_{1}\rangle=r|1,0\rangle+t|0,1\rangle$, producing $\Pi_{2}$. The unused port defines the inconclusive outcome $\Pi_{0}$, monitored at $D_{0}$.
@@
For each value of $|t|$, the idler counts $(N_{1},N_{2},N_{0})$ determine $P_{S}=(N_{1}+N_{2})/N_{\mathrm{tot}}$ and $P_{I}=N_{0}/N_{\mathrm{tot}}$, while the signal visibility $V$ is extracted from single-photon fringes. In the ideal regime of pure idler markers, IDP theory predicts $P_{I}=|\langle\phi_{1}|\phi_{2}\rangle|=|t|$ and $P_{S}=1-|t|$, while the ZWM analysis gives $V=|t|$ and $D=\sqrt{1-|t|^{2}}$. The experiment therefore directly tests the operational identifications $P_{I}=V$ and $D^{2}+V^{2}=1$. Details of optical layout, mode matching, loss, and detector efficiency are provided in Appendix F.
IV Mixed idler marker states and thermal noise
@@
In realistic implementations, the idler mode that propagates between the two SPDC crystals is not perfectly isolated and may interact with environmental degrees of freedom such as thermal background radiation and loss channels [giri1, boyd]. This coupling renders the conditional idler markers associated with emission from crystals A and B mixed states, denoted by $\rho_{A}$ and $\rho_{B}$, rather than pure states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$. As a result, the interference visibility in the signal arm is no longer determined solely by the inner product of pure marker states, but is instead upper bounded by the Uhlmann fidelity $F(\rho_{A},\rho_{B})$ [Uhlmann76] between the two mixed markers; see Appendix G. By a standard purification argument one finds
@@
which reduces to the ideal relation $V=|\langle\phi_{1}|\phi_{2}\rangle|$ in the absence of thermal noise.
From the theory of unambiguous quantum state discrimination (IDP), the optimal average failure probability $P_{I}^{\mathrm{opt}}$ for distinguishing $\rho_{A}$ from $\rho_{B}$ with equal priors satisfies
$P_{I}^{\mathrm{opt}}\geq F(\rho_{A},\rho_{B}),$ (17)
@@
$V\;\leq\;F(\rho_{A},\rho_{B})\;\leq\;P_{I}^{\mathrm{opt}}.$ (18)
Thus, the presence of thermal photons reduces the observed signal single-photon fringe visibility by increasing the distinguishability of the mixed conditional idler markers [giri1], while the IDP discrimination probability provides the tight operational upper bound. In the zero-noise limit these inequalities collapse to
$V=F(\rho_{A},\rho_{B})=P_{I}^{\mathrm{opt}},$ (19)
@@
Physical intuition: The situation is closely analogous to the double-slit experiment: interference occurs only when it is impossible to know which path the photon took. In the ZWM interferometer, the two crystals play the role of the two slits, and the idler photon acts as the which-path marker. If we ignore the idler, the signal photon shows high-visibility interference. If we measure the idler strongly to learn which crystal emitted the pair, interference disappears. The key new ingredient in our proposal is the use of the Ivanovic–Dieks–Peres (IDP) unambiguous discrimination measurement, which has three possible outcomes: it sometimes identifies the emitting crystal (destroying interference) but sometimes returns an inconclusive result, in which case interference remains perfectly intact. Thus, instead of switching interference fully on or off, IDP allows us to separate the data into wave-like and particle-like subsets, revealing a direct operational meaning of visibility as the unavoidable failure probability in zero-error state discrimination. Visibility quantifies the optimal failure probability in zero-error discrimination, whereas distinguishability quantifies the optimal hypothesis-test success probability.
V Conclusion
@@
Conclusion.— We have shown that in induced-coherence interferometry the standard complementarity relation $D^{2}+V^{2}=1$ acquires a direct operational interpretation: for pure which-crystal idler markers in the low-gain ZWM interferometer, the singles visibility equals the optimal inconclusive probability of the Ivanovic–Dieks–Peres unambiguous discrimination measurement, $V=P_{I}^{\mathrm{opt}}$. Thus wave–particle duality expresses the optimality boundary of a zero-error hypothesis test on the idler mode, demonstrating that accessible which-path information is fundamentally measurement-limited rather than interferometer-limited.
Comparison with the coincidence-based extraction of Torres et al. shows that standard correlation measurements underestimate distinguishability for partially overlapping marker states ($0<T<1$), while both approaches yield $D=0$ when the markers become identical. With thermal noise, we derived the hierarchy $V\leq F(\rho_{A},\rho_{B})\leq P_{I}^{\mathrm{opt}}$, linking visibility to Uhlmann fidelity and fidelity to optimal discrimination.
The IDP measurement naturally introduces a vacuum-flagged inconclusive outcome that functions analogously to an erasure syndrome in quantum error syndrome [Bennett1996], suggesting that control of interference can be governed by measurement design rather than physical indistinguishability [Aqua]. These results motivate discrimination-optimized induced-coherence imaging and sensing approaches [Lloyd2008, Shapiro2015, giri1], and open a route toward experiments demonstrating USD-controlled visibility.
Appendix A Schrödinger-picture derivation of the ZWM two-crystal state
In the low-gain regime of SPDC [zou1991induced], the interaction Hamiltonian is
@@ Appendix A Schrödinger-picture derivation of the ZWM two-crystal state
and the corresponding weak unitary evolution generates the single-crystal state
$|\Psi_{\mathrm{single}}\rangle\approx|0\rangle+g\,|1_{S},1_{I}\rangle,$ (21)
where $g$ is a small, dimensionless gain parameter proportional to the nonlinear coupling $\kappa$ (low-gain condition $|g|\ll 1$). For two coherently pumped crystals A and B,
$\begin{split}|\psi_{SI}\rangle=&\frac{1}{\sqrt{2}}\,|1_{S_{1}},0_{S_{2}}\rangle\Big(r\,|1_{I_{2}},0_{I_{3}}\rangle+t\,|0_{I_{2}},1_{I_{3}}\rangle\Big)\\ &+\frac{1}{\sqrt{2}}\,|0_{S_{1}},1_{S_{2}}\rangle\,|0_{I_{2}},1_{I_{3}}\rangle.\end{split}$ (22)
The first term represents the creation of a signal–idler photon pair in the first nonlinear crystal, whereas the second term corresponds to pair generation in the second crystal.
A.1 Conditional idler marker states
We now define the which-crystal alternatives as [Torres2024]:
  • Crystal A emits a pair: signal state $|1_{S_{1}},0_{S_{2}}\rangle$;
  • Crystal B emits a pair: signal state $|0_{S_{1}},1_{S_{2}}\rangle$.
Conditioned on these alternatives, the corresponding idler states are:
Emission from crystal A.
Projecting the two-crystal state of Eq. (22) onto the signal branch $|1_{S_{1}},0_{S_{2}}\rangle$ isolates the contribution from crystal A. Up to normalization, the corresponding idler state is
$|\phi_{1}\rangle=r\,|1,0\rangle_{I_{2},I_{3}}+t\,|0,1\rangle_{I_{2},I_{3}}.$ (23)
where $r$ and $t$ denote the effective reflection and transmission amplitudes for the idler mode, incorporating both the pump strengths and the interferometric phase accumulation.
Emission from crystal B.
Project onto the signal branch $|0_{S_{1}},1_{S_{2}}\rangle$. The idler state is then
$|\phi_{2}\rangle=|0,1\rangle_{I_{2},I_{3}}.$ (24)
These two pure idler states encode which-crystal information. Their overlap is
$\langle\phi_{1}|\phi_{2}\rangle=r^{*}\langle 1,0|0,1\rangle+t^{*}\langle 0,1|0,1\rangle=t^{*},$ (25)
@@ A.1 Conditional idler marker states
$|\langle\phi_{1}|\phi_{2}\rangle|=|t|.$ (26)
A.2 Signal visibility and distinguishability
To obtain the visibility in the signal arm, we trace over the idler and recombine the signal modes $S_{1}$ on a balanced beam splitter. If both crystals are pumped [Lahirinon, Krenn2022, Shafiee2024] by the same laser with equal intensities and identical phase-matching and coupling conditions Eqs.21, the amplitudes for pair creation are $g_{1}\quad\text{and}\quad g_{2}$, and under symmetric pumping one has
$|g_{1}|=|g_{2}|.$
Conditioned on the event that a pair is generated, the a priori probabilities that the pair originated from crystal A or B are proportional to the squared amplitudes:
$\eta_{1}=\frac{|g_{1}|^{2}}{|g_{1}|^{2}+|g_{2}|^{2}},\qquad\eta_{2}=\frac{|g_{2}|^{2}}{|g_{1}|^{2}+|g_{2}|^{2}}.$ (27)
If $|g_{1}|=|g_{2}|$, then
$\eta_{1}=\eta_{2}=\frac{1}{2}.$ (28)
In particular, for equal priors $\eta_{1}=\eta_{2}=\tfrac{1}{2}$, the fringe visibility $V$ is
$V=|\langle\phi_{1}|\phi_{2}\rangle|=|t|.$ (29)
A natural measure of which-path distinguishability for two pure alternatives with equal priors is
$D=\sqrt{1-|\langle\phi_{1}|\phi_{2}\rangle|^{2}}=\sqrt{1-|t|^{2}}=|r|.$ (30)
Combining Eqs. (29) and (30), we obtain the standard complementarity relation
$D^{2}+V^{2}=(1-|t|^{2})+|t|^{2}=1.$ (31)
A.3 Imperfect mode overlap (optional extension)
@@
Appendix B Visibility and IDP failure probability for pure markers
In the low-gain ZWM configuration, the two which-crystal idler marker states are
@@ Appendix B Visibility and IDP failure probability for pure markers
with $|r|^{2}+|t|^{2}=1$, defined on the two-dimensional single-photon subspace of the idler output modes. Their overlap is
$|.s:=|\innerproduct{\phi_{1}}{\phi_{2}}|=|t$|. (36)
Tracing out the idler and recombining the signal modes at a balanced beam splitter yields single-photon interference fringes with visibility
[…]
and a natural distinguishability measure
$|,D=\sqrt{1-|\innerproduct{\phi_{1}}{\phi_{2}}|^{2}}=\sqrt{1-s^{2}}=|r$|, (38)
so that $D^{2}+V^{2}=1$.
@@
The optimal Ivanovic–Dieks–Peres (IDP) strategy [Ivanovic, Dieks, Peres] minimizes the inconclusive probability
@@
over all such POVMs. For two pure states with equal priors it is well known that the minimal value is
@@
with the corresponding success probability $P_{S}^{\mathrm{opt}}=1-P_{I}^{\mathrm{opt}}$. For completeness, an explicit construction of the optimal POVM for the ZWM markers is given in Appendix A.
Combining Eqs. (37) and (41) we obtain
$V=P_{I}^{\mathrm{opt}},$ (42)
so that the usual complementarity relation can be written as
$D^{2}+\bigl(P_{I}^{\mathrm{opt}}\bigr)^{2}=1.$ (43)
Thus, for pure which-path markers in a balanced two-path interferometer, wave–particle duality is equivalent to the optimality condition of unambiguous quantum-state discrimination.
Appendix C Optimal IDP discrimination for the ZWM idler states
We now derive the optimal Ivanovic–Dieks–Peres (IDP) measurement [Clarke2001IDP, Danan2013Trajectories, Herzog2005MixedUSD, mohseni] for the ZWM idler states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$ defined in Eqs. (23) and (24), and we show explicitly that the optimal inconclusive probability equals the ZWM visibility.
C.1 Orthogonal complements of the idler states

### v3

@@ VII Experimental Protocol
The IDP measurement can be implemented [Barnett, mohseni] directly on the idler modes of a two-crystal ZWM interferometer [Torres2024]. Two coherently pumped SPDC crystals (A,B) generate signal–idler pairs in the low-gain regime. The idler from A traverses an object $S$ with transmissivity $(t,r)$, while the idler from B bypasses it. The signal modes are combined at a balanced beam splitter to measure the single-photon visibility $V$.
@@
To realize the IDP POVM $\{\Pi_{inc},\Pi_{A},\Pi_{B}\}$ on the idler as shown in Fig. 1. Detector $D_{1}$ implements the projection onto $|1,0\rangle$, yielding conclusive outcome $\Pi_{A}$. Detector $D_{2}$ receives the output of a variable beam splitter that interferes the idler modes with a calibrated phase such that the detected mode is $t^{*}|1,0\rangle-r^{*}|0,1\rangle$, orthogonal to $|\phi_{A}\rangle=r|1,0\rangle+t|0,1\rangle$, producing $\Pi_{B}$. The unused port defines the inconclusive outcome $\Pi_{inc}$, monitored at $D_{0}$.
@@
For each value of $|t|$, the idler counts $(N_{1},N_{2},N_{0})$ determine $P_{suc}=(N_{1}+N_{2})/N_{\mathrm{tot}}$ and $P_{inc}=N_{0}/N_{\mathrm{tot}}$, while the signal visibility $V$ is extracted from single-photon fringes. In the ideal regime of pure idler markers, IDP theory predicts $P_{inc}=|\langle\phi_{A}|\phi_{B}\rangle|=|t|$ and $P_{suc}=1-|t|$, while the ZWM analysis gives $V=|t|$ and $D=\sqrt{1-|t|^{2}}$. The experiment therefore directly tests the operational identifications $P_{inc}=V$ and $D^{2}+V^{2}=1$. Details of optical layout, mode matching, loss, and detector efficiency are provided in Appendix H.
VIII Mixed idler marker states and thermal noise
@@
In realistic implementations, the idler $\hat{a}_{3}^{1}$ mode that propagates between the two SPDC crystals is not perfectly isolated. The object $S$ placed between the two crystal. Object interact with environmental degrees of freedom such as mean thermal photon $\langle\hat{a}_{4}^{\dagger}\hat{a}_{4}\rangle=N_{B}$ and loss channels [giri1, boyd]. This coupling renders the conditional idler markers $I$ associated with emission from crystals A and B mixed states. The loss channel denoted by $\rho_{A}$ and $\rho_{B}$, rather than pure states $|\phi_{A}\rangle$ and $|\phi_{B}\rangle$. As a result, the interference visibility in the signal arm is no longer determined solely by the inner product of pure marker states. It instead upper bounded by the Uhlmann fidelity $F(\rho_{A},\rho_{B})$ [Uhlmann76] between the two mixed markers; see Appendix I. By a standard purification argument one finds
@@
which reduces to the ideal relation $V=|\langle\phi_{A}|\phi_{B}\rangle|$ in the absence of thermal noise.
From the theory of unambiguous quantum state discrimination (IDP), the optimal average failure probability $P_{inc}^{\mathrm{opt}}$ for distinguishing $\rho_{A}$ from $\rho_{B}$ with equal priors satisfies
$P_{inc}^{\mathrm{opt}}\geq F(\rho_{A},\rho_{B}),$ (51)
@@
$V\;\leq\;F(\rho_{A},\rho_{B})\;\leq\;P_{inc}^{\mathrm{opt}}.$ (52)
Thus, the presence of thermal photons reduces the observed signal single-photon fringe visibility by increasing the distinguishability of the mixed conditional idler markers [giri1]. The IDP discrimination probability provides the tight operational upper bound. In the zero-noise limit these inequalities collapse to
$V=F(\rho_{A},\rho_{B})=P_{inc}^{\mathrm{opt}},$ (53)
@@
Physical intuition: The situation is closely analogous to the double-slit experiment: interference occurs only when it is impossible to know which path the photon took. In the ZWM interferometer, the two crystals play the role of the two slits, and the idler photon acts as the which-path marker. If we ignore the idler, the signal photon shows high-visibility interference. If we measure the idler strongly to learn which crystal emitted the pair, interference disappears. The key new ingredient in our proposal is the use of the Ivanovic–Dieks–Peres (IDP) unambiguous discrimination measurement. Which has three possible outcomes: it sometimes identifies the emitting crystal (destroying interference) but sometimes returns an inconclusive result. In which case interference remains perfectly intact. Thus, instead of switching interference fully on or off, IDP allows us to separate the data into wave-like and particle-like subsets. Revealing a direct operational meaning of visibility as the unavoidable failure probability in zero-error state discrimination. Visibility quantifies the optimal failure probability in zero-error discrimination, whereas distinguishability quantifies the optimal hypothesis-test success probability.
In table 1, operational interpretation of visibility $V$ and distinguishability $D$ under different quantum decision strategies for which-crystal discrimination in induced-coherence interferometry.
IX Conclusion
@@
We have shown that in induced-coherence interferometry the standard complementarity relation $D^{2}+V^{2}=1$ admits a direct operational interpretation. In the low-gain ZWM interferometer with pure which-crystal idler markers. The singles visibility equals the optimal inconclusive probability of Ivanovic–Dieks–Peres unambiguous state discrimination, $V=P_{inc}^{\mathrm{opt}}$. Wave–particle duality therefore corresponds to the optimality boundary of a zero-error quantum hypothesis test on the idler. Demonstrating that accessible which-crystal information is fundamentally measurement-limited rather than interferometer-limited.
Complementarily, we showed that the distinguishability $D_{Hel}$ coincides with the Helstrom trace-distance distinguishability. The Helstrom bound yields the optimal minimum-error probability for identifying which crystal emitted the photon pair in a single-shot idler measurement. In this sense, visibility and distinguishability acquire dual operational meanings as optimal failure and success probabilities of quantum decision strategies.
Compared with coincidence-based extraction methods, which provide only a proxy for which-crystal information. Optimal discrimination reveals greater distinguishability for partially overlapping marker states. In the presence of thermal noise, this picture generalizes to the hierarchy $V\leq F(\rho_{A},\rho_{B})\leq P_{inc}^{\mathrm{opt}}$, linking visibility, Uhlmann fidelity, and optimal discrimination.
These results motivate discrimination-optimized induced-coherence imaging and sensing approaches [Lloyd2008, Shapiro2015, giri1], and open a route toward experiments demonstrating USD-controlled visibility. Beyond fundamental interest, induced coherence is crucial in quantum imaging with undetected photons [Lemos2014], induced-coherence imaging theory [Lahiri2019], and phase-sensitive sensing [Shapiro2015].
Acknowledgements.
L. T. thanks David Vitali (UNICAM) for helpful comments on the manuscript. L. T. also acknowledges C. M. Chandrashekar (IISc) for financial support.
Appendix A Schrödinger-picture derivation of the ZWM state
The interaction Hamiltonian for a single SPDC crystal is
@@ Appendix A Schrödinger-picture derivation of the ZWM state
with weak unitary evolution
[…]
To first order in the small gain parameter $g\propto\kappa\Delta t$, one obtains
[…]
For two coherently pumped crystals $A$ and $B$ we write the joint state (up to a global vacuum component)
[…]
where $I_{A}$ and $I_{B}$ are the idler modes associated with crystals $A$ and $B$. Alignment of the idler paths between the crystals makes these modes physically identical, $I_{A}\equiv I_{B}\equiv I$, so the single-pair component can be written as
[…]
An object $S$ placed in the idler path is modelled as a lossless beam splitter with amplitudes $t$ and $r$ acting on the idler mode and an auxiliary vacuum mode $V$:
$\hat{a}_{I}^{\dagger}$ $\to r\hat{a}_{I_{2}}^{\dagger}+t\hat{a}_{I_{3}}^{\dagger},$ (59)
$\hat{a}_{V}^{\dagger}$ $\to-t^{*}\hat{a}_{I_{2}}^{\dagger}+r^{*}\hat{a}_{I_{3}}^{\dagger},$ (60)
@@
[…]
Three regimes.
The induced-coherence visibility is controlled by the overlap of the conditional idler marker states.
(i) Transparent object ($t=1$, $r=0$): both emission alternatives prepare the same idler state, $|\phi_{A}\rangle=|\phi_{B}\rangle$, hence $\langle\phi_{A}|\phi_{B}\rangle=1$ and the reduced signal state retains full coherence. This yields maximal single-photon visibility $V=1$.
(ii) Opaque object ($t=0$, $r=1$): the idler associated with emission from crystal A exits in a mode orthogonal to the idler of crystal B, so that $\langle\phi_{A}|\phi_{B}\rangle=0$. Tracing over the idler removes all signal-path coherence, and the singles interference vanishes ($V=0$).
(iii) Intermediate transmissivity ($0<|t|<1$): the marker states are nonorthogonal, $0<|\langle\phi_{A}|\phi_{B}\rangle|<1$, so the reduced signal density matrix retains a partial off-diagonal coherence proportional to $\langle\phi_{A}|\phi_{B}\rangle$. The visibility is
$V=|\langle\phi_{A}|\phi_{B}\rangle|=|t|.$ (62)
In the low gain limit, conditioned on emission from crystal $A$, the idler state after $S$ is therefore
[…]
while emission from $B$ corresponds to an idler bypass that prepares
[…]
These are the marker states used in the main text.
Tracing over the idler and recombining $S_{1}$ and $S_{2}$ on a balanced beam splitter yields a reduced signal density matrix
[…]
from which the fringe visibility follows as
$|.V=\frac{2|\rho_{12}|}{\rho_{11}+\rho_{22}}=\absolutevalue{\langle\phi_{1}|\phi_{2}\rangle}$. (66)
Appendix B Imperfect mode overlap
@@
Appendix C Helstrom minimum–error discrimination for the ZWM idler marker states
The minimum–error quantum hypothesis testing (Helstrom bound) to the induced-coherence (ZWM) idler-marker states. The task is to infer which crystal fired from a single-shot measurement on the idler mode.
C.0.1 Binary hypothesis test, POVM, and error probabilities
We consider a binary quantum hypothesis test [Helstrom] with priors $p_{0},p_{1}\geq 0$ and $p_{0}+p_{1}=1$:
$H_{0}:\ \hat{\rho}=\hat{\rho}_{0},\ p=p_{0},\qquad H_{1}:\ \hat{\rho}=\hat{\rho}_{1},\ p=p_{1}.$ (70)
Bob performs a dichotomous POVM $\{\Pi_{0},\Pi_{1}\}$ with outcomes $\mu\in\{0,1\}$ satisfying
@@ C.0.1 Binary hypothesis test, POVM, and error probabilities
The decision rule is: decide $H_{0}$ when $\mu=0$ and decide $H_{1}$ when $\mu=1$.
The probability errors are, respectively,
@@
The Bayes average error probability is
@@
C.0.2 ZWM idler marker states and density operators
In the low-gain induced-coherence (ZWM) regime [Torres2024], the idler marker lives in the single-photon two-mode subspace spanned by $\{|1,0\rangle,\ |0,1\rangle\}$. The two marker kets are
with
$|r|^{2}+|t|^{2}=1,\qquad s:=|\langle\phi_{1}|\phi_{2}\rangle|=|t|.$ (76)
We map these to the hypotheses
$H_{0}:\ \hat{\rho}_{0}=|\phi_{1}\rangle\langle\phi_{1}|,\qquad H_{1}:\ \hat{\rho}_{1}=|\phi_{2}\rangle\langle\phi_{2}|.$ (77)
In the ordered basis $\{|1,0\rangle,\ |0,1\rangle\}$, the density operators read
$\hat{\rho}_{0}=\begin{pmatrix}|r|^{2}&rt^{*}\\ r^{*}t&|t|^{2}\end{pmatrix},\qquad\hat{\rho}_{1}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$. (78)
C.0.3 Helstrom matrix and optimal POVM (general priors)
The Helstrom matrix is defined as
$\gamma:=p_{0}\hat{\rho}_{0}-p_{1}\hat{\rho}_{1}.$ (79)
Let the spectral decomposition of $\gamma$ be
$\gamma=\sum_{\mu}\gamma_{\mu}|\mu\rangle\langle\mu|.$ (80)
The Helstrom POVM is obtained by projecting onto the positive part $\gamma_{+}$ of $\gamma$:
@@ C.0.3 Helstrom matrix and optimal POVM (general priors)
The minimum achievable Bayes error is
[…]
C.0.4 Equal priors $p_{0}=p_{1}=\tfrac{1}{2}$: explicit ZWM solution
From here onward we assume equiprobable hypotheses,
$p_{0}=p_{1}=\frac{1}{2},\qquad\gamma=\frac{1}{2}(\hat{\rho}_{0}-\hat{\rho}_{1}).$ (83)
Using (78) we obtain
$\gamma=\frac{1}{2}\begin{pmatrix}|r|^{2}&rt^{*}\\ r^{*}t&|t|^{2}-1\end{pmatrix}=\frac{1}{2}\begin{pmatrix}|r|^{2}&rt^{*}\\ r^{*}t&-|r|^{2}\end{pmatrix}$. (84)
Define
$a:=|r|^{2},\qquad b:=rt^{*},$ (85)
so that
$\gamma=\frac{1}{2}\begin{pmatrix}a&b\\ b^{*}&-a\end{pmatrix}$. (86)
Eigenvalues.
[…]
$0$ $=\det\!\left(\frac{1}{2}\begin{pmatrix}a&b\\ b^{*}&-a\end{pmatrix}-\lambda\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right)=\det\begin{pmatrix}\frac{a}{2}-\lambda&\frac{b}{2}\\[2.0pt] \frac{b^{*}}{2}&-\frac{a}{2}-\lambda\end{pmatrix}$
$=\left(\frac{a}{2}-\lambda\right)\left(-\frac{a}{2}-\lambda\right)-\frac{|b|^{2}}{4}=\lambda^{2}-\frac{1}{4}\bigl(a^{2}+|b|^{2}\bigr).$ (87)
Hence
$\lambda_{\pm}=\pm\frac{1}{2}\sqrt{a^{2}+|b|^{2}}.$ (88)
For the ZWM parameters,
$a^{2}+|b|^{2}=|r|^{4}+|r|^{2}|t|^{2}=|r|^{2}\bigl(|r|^{2}+|t|^{2}\bigr)=|r|^{2},$ (89)
so
$\boxed{\lambda_{\pm}=\pm\frac{|r|}{2}.}$ (90)
Trace norm.
Since $\gamma$ is Hermitian, $\|\gamma\|_{1}=\sum_{\pm}|\lambda_{\pm}|$, hence
$\boxed{\|\gamma\|_{1}=\left|\frac{|r|}{2}\right|+\left|\!-\frac{|r|}{2}\right|=|r|.}$ (91)
Helstrom minimum error and maximum success.
Inserting (91) into (82) yields
$\boxed{P_{\rm err}^{\min}=\frac{1}{2}\bigl(1-|r|\bigr),\qquad P_{\rm s}^{\max}=1-P_{\rm err}^{\min}=\frac{1}{2}\bigl(1+|r|\bigr).}$ (92)
Using the overlap $s=|\langle\phi_{1}|\phi_{2}\rangle|=|t|$ and $|r|=\sqrt{1-|t|^{2}}=\sqrt{1-s^{2}}$, we may rewrite
$\boxed{P_{\rm err}^{\min}=\frac{1}{2}\Bigl(1-\sqrt{1-s^{2}}\Bigr)=\frac{1}{2}\Bigl(1-\sqrt{1-|t|^{2}}\Bigr).}$ (93)
C.0.5 Explicit Helstrom POVM for the ZWM markers
By definition, the Helstrom measurement projects onto the eigenvectors of $\gamma$:
$\Pi_{0}=|\omega_{+}\rangle\langle\omega_{+}|,\qquad\Pi_{1}=|\omega_{-}\rangle\langle\omega_{-}|,$ (94)
where
$\gamma|\omega_{\pm}\rangle=\lambda_{\pm}|\omega_{\pm}\rangle.$ (95)
Write the complex amplitudes as
$r=|r|e^{i\alpha},\qquad t=|t|e^{i\beta},\qquad\varphi:=\arg(rt^{*})=\alpha-\beta,$ (96)
and define an angle $\theta\in[0,\pi/2]$ by
$\cos\theta=|r|,\qquad\sin\theta=|t|=s.$ (97)
A convenient normalized eigenbasis of $\gamma$ is then
$\boxed{|\omega_{+}\rangle=\cos\frac{\theta}{2}\,|1,0\rangle+e^{-i\varphi}\sin\frac{\theta}{2}\,|0,1\rangle,}$ (98)
$\boxed{|\omega_{-}\rangle=\sin\frac{\theta}{2}\,|1,0\rangle-e^{-i\varphi}\cos\frac{\theta}{2}\,|0,1\rangle.}$ (99)
With (94) this fully specifies the optimal Helstrom POVM for the ZWM idler markers.
Appendix D Connection to visibility and operational complementarity
Helstrom distinguishability, assuming equal priors $p_{0}=p_{1}=\tfrac{1}{2}$, define the Helstrom operator
$|.\gamma=\frac{1}{2}(\rho_{1}-\rho_{2}),\qquad\rho_{j}=\ket{\phi_{j}}\bra{\phi_{j}}$. (100)
The Helstrom (trace-distance) distinguishability is
$D_{\mathrm{Hel}}\equiv\|\gamma\|_{1}=\frac{1}{2}\|\rho_{1}-\rho_{2}\|_{1}.$ (101)
For two pure states, the trace distance has the closed-form identity
[…]
Using Eq. (66), this yields
$D_{\mathrm{Hel}}=\sqrt{1-V^{2}}.$ (103)
D.0.1 Helstrom form of the duality relation
Squaring Eq. (103) and adding $V^{2}$ immediately gives
$\boxed{D_{\mathrm{Hel}}^{2}+V^{2}=1,}$ (104)
which coincides exactly with the standard pure-state duality relation Eq. (8). Thus, in the ZWM interferometer with pure idler markers, the “particle-like” distinguishability $D$ appearing in the duality relation is precisely the Helstrom trace-distance distinguishability for the minimum-error “which-crystal fired?” discrimination task.
D.0.2 Connection to the Helstrom bound
The minimum achievable average error probability for this task is
$P_{\mathrm{err}}^{\min}=\frac{1}{2}\bigl(1-\|\gamma\|_{1}\bigr)=\frac{1}{2}\bigl(1-D_{\mathrm{Hel}}\bigr)=\frac{1}{2}\Bigl(1-\sqrt{1-V^{2}}\Bigr).$ (105)
For the ZWM parameterization $|r|^{2}+|t|^{2}=1$, one has
$D_{\mathrm{Hel}}=|r|,\qquad V=|t|,$ (106)
so that
$|r|^{2}+|t|^{2}=1\quad\Longleftrightarrow\quad D_{\mathrm{Hel}}^{2}+V^{2}=1,\qquad P_{\mathrm{err}}^{\min}=\frac{1}{2}(1-|r|).$ (107)
Appendix E Optimal IDP discrimination for the ZWM idler states
We now derive the optimal Ivanovic–Dieks–Peres (IDP) measurement [Clarke2001IDP, Danan2013Trajectories, Herzog2005MixedUSD, mohseni] for the ZWM idler states $|\phi_{1}\rangle$ and $|\phi_{2}\rangle$, we show explicitly that the optimal inconclusive probability equals the ZWM visibility.
E.1 Orthogonal complements of the idler states